OpenStudy (anonymous):

Is this correct? Use the even-odd properties of the trig functions to find the exact value of the given function. sin(-pi/4) = - (sqrt2)/2

OpenStudy (mertsj):

yep

myininaya (myininaya):

\[\sin(\frac{-\pi}{4})=-\sin(\frac{\pi}{4})\]

myininaya (myininaya):

by using the fact that sin(x) is odd

OpenStudy (anonymous):

thanks myininaya, can you please help me with another question i have posted?